Run Coordination Meetings: Spring 2016 Run

From GlueXWiki
Jump to: navigation, search

Meeting General Agenda

Connect Information

To connect from the outside by phone please do the following

1.) To join via a Web Browser, go to the page [1] https://bluejeans.com/660743227.

2.) To join via Polycom room system go to the IP Address: 199.48.152.152 (bjn.vc) and enter the meeting ID: 660743227.

3.) To join via phone, use one of the following numbers and the Conference ID: 660743227.

4.) The moderator code is 8394

Dial Either +1 408 740 7256 or +1 888 240 2560 for US or Canada

4.) More information on connecting to bluejeans is available. The old ESNET and SeeVough Information is archived here.

Specific instructions for connecting from JLab's Hall D conference room:

  • Turn polycon on if necessary (do it before turning the computer on)
  • With the polycon, place a call at 199.48.152.152
  • Press # to enable the polycom keypad, then enter the meeting id: 660743227 and #
  • You may have to unmute the microphone: #*4
  • Turn the computer on if needed

Beam and Hall configuration

  • Beam energy expected: 12.047 GeV (1.090 GeV/linac and 0.123 GeV for the injector and -0.066 GeV of synchrotron radiation losses.
  • Solenoid on at 1200A. Any field changes should be avoided if possible.
  • Beam current: CW 5 na-3 μA. 250 MHz frequency. Default current and radiator: 100 nA/1.12·10-4 R.L.
  • Radiators:
    • Goniometer diamond radiators: 3 test diamonds
      • J1a50 (50 μm, 4.1·10-4 R.L.),
      • JD70-118 (20 μm, 1.6·10-4 R.L.), middot;10-4 R.L.),
      • JD70-119 (20 μm, 1.6·10-4 R.L.), middot;10-4 R.L.),
    • and 3 test foils (two 10 μm Al foils, and one 100 μm Al foil with a holed 10 μm Al foil pressed on it (3mm hole)).
    • Amorphous
      • 1.5 µm Al (1.7·10-5 R.L.)
      • 10 µm Al (11.2·10-5 R.L.)
      • 30 µm Al (33.7·10-5 R.L.)
      • Possibly a ladder with thiner diamonds (20 μm diamonds) will be installed during second half of the run. To be confirmed.
  • Collimator hole: 5 mm diameter and possibly 3.4 mm diameter. We may use a plug to reduce the 5mm hole to 1.05 mm in order to commission and use the TAC.
  • Target: LH2 target: cryogenic target

General Information

This document describes the run plan for the Hall D Spring 2016 run.

  • Schedule for the Spring 2016 run. (Dates are according to the long term calendar.)
    1. Jan. 28th-Feb. 3rd: Electron beam restoration
    2. Feb. 4th-Apr. 14th: Hall D Spring run.
    3. Every Tuesdays: 12h of beam study. Every Thursdays, if necessary: 8h of RF recovery.
  • Goals for the Spring 2016 run.
  1. Beam transport and control commissioning (accelerator responsibility):
    1. Position Fast Feedback: Trent Allison, Brian Bevins. Expected time: 40h.
    2. nA BPM: Trent Allison, Brian Bevins. 8h.
    3. Beam line model test: Todd Satogata). Expected time: 2 h.
    4. Test procedure for fast Active Collimator calibration
  2. Continue commissioning of the Hall D equipment (Hall D responsibility):
    1. Trigger: (Alex S).
    2. DAQ (Sergey F).
    3. CDC high voltage scan (Ben. Z.).
    4. Goniometer commissioning at 12 GeV (Hovanes/Paul/Ken)
    5. π0 calibration. (Elton S.)
    6. Trigger Level-3 commissioning. (David L.)
    7. Establish absolute photon flux with PS (Sacha) and with TAC (Hovanes)
    8. Check the effect of the quadrupole on the microscope data (Alex. Barnes).
    9. 3.4 mm hole run if polarized beam is established with thin diamonds.
    10. (TRD, in parallel with other tasks)
    11. (Polarimeter (Nathan S.), in parallel with other tasks)


Runplan

  1. Solenoid tests
    • Solenoid tests have priority.
    • No beam required. Will start as soon as the magnet is cold. May require a long access (at least 8h, day) if the cooling is not finished by the start of the run.
  2. Re-establish electron and photon beams (assume 3 days).
    • Ion Chambers trip threshold must be set.
    • Tune electron beam with radiator retracted
      • Amorphous Radiator (and diamond) must be retracted. Collimator should be in blocking position. CDC/FDC are off
      • Radiation levels should be closely monitored by MCC crew.
        • Typical levels are (for radiator retracted, collimator fully blocking, electron beam current 50 nA)
          • RAD102_P1 ~ 3 mrad/hr (tagger area, gammas, between tagger and dump)
          • RAD102_P2 ~ 15 mrad/hr (tagger area, gammas, near electronics racks).
          • RAD102_P3 ~ 0.1 mrem/hr (tagger area, neutrons, near electronics racks)
          • RAD508_P1 < 0.1 mrad/hr (collimator cave, gammas).
          • RAD508_P2 < 0. mrad/hr (Hall. Neutrons).
      • Hall D tagger CARMS, Ion Chamber and BLM threshold for beam trips should not be changed without approval of Hall D leader.
    • Establish good photon beam
      • Insert 10-5 RL radiator. (Call MCC and ask to mask the FSD each time the radiator is moved)
        • Typical levels are (for 2*10^-5 radiator, electron beam current 50 nA)
          • RAD102_P1= 2-10 mrad/hr (tagger area, gammas, between tagger and dump)
          • RAD102_P2= 10-20 mrad/hr (tagger area, gammas, near electronics racks)
          • RAD102_P3= ~0.5 mrem/hr (tagger area, neutrons, near electronics racks)
          • RAD508_P1= ~10 mrad/hr (collimator cave, gammas).
          • RAD508_P2 < ~ 0. mrad/hr (Hall. Neutrons).
    • Once radiation levels are good, proceed with radiation level study (1h):
      • Run 5 min without radiator (Call MCC and ask to mask the FSD each time the radiator is moved). Beam current: 50nA. Repeat at 100 nA and 200 nA
      • Run 5 min with the 10^-4 radiator (Call MCC). Beam current: 50nA. Repeat at 100 nA and 200 nA
      • Run 3 min with the 3*10^-4 radiator (Call MCC). Beam current: 50nA. Repeat at 100 nA and 200 nA
    • Active Collimator calibration.
      • re-Insert the 2*10^-5 radiator (Call MCC).
      • Determine maximum photon beam transmission by doing 2D scans (x: beam scan, y: beam scan. ±5mm with 0.5mm steps). Stay 5 seconds with stable beam for each steps Remind MCC that they should use only the last corrector (5C11A) to adjust the beam position. If the adjustment results in unacceptable e- beam position in the tagger dump, remind MCC that they should use the tagger magnet only to readjust the e- beam position in the tagger dump.Meanwhile, monitor the rate at the active target and rates (including coinc.) in the pair spectrometer.
      • Record on the white board what is the profiler position corresponding to the center of the A.C.
    • Verify with MCC that current/position/energy locks are on, if possible.
    • before Fast Feedback: Quick test turning the Tagger quadrupole on to make sure the beam is centered (no quad. steering). (30min)
      • Current setting (MCC controls it): 3.2968A. This yields a 520 G/cm gradient. Ask MCC to make sure quad. current polarity was properly set.
      • Start a MyaViewer strip chart with quad current and AD00c x and y positions (tagger dump position).
      • Once quad. is on, check that it did not steer the beam on AD00c. Repeat quad on/off and check correlation a few times.
      • If beam is steered by the quad. more than ~1mm, we will need an opportunistic access to mechanically center it (Todd Satogata), and then repeat the test.
  3. Accelerator Fast Feedback commissioning (10×4h), nA BPM commissioning (2×4h) Optic test (2h), Active Collimator calibration test (2h), and Hall D commissioning
    • Solenoid tests (cont.), priority.
    • FFB commissioning will take ~4h/days for a week to 10 days.
      • Beam need to be well tuned (no tail) in order to allow enough excursion for testing the feedback system, without tripping the ion chambers.
    • FOPT optic test (2h).
    • HV scans for CDC (Beni, 4×1h)
      • Need stable beam condition
      • Solenoid must be on.
      • 4 CDC voltage setting: 2100V, 2125V, 2050V and 2000V.
    • Hall D trigger test (part of it can be done in parallel with other tasks, including FFB/nA/FOPT work. Alex. S. 2h/day except Feb. 10th-15th)
      • Production trigger optimization (FCal&BCal, PS, FCal/BCal) 5×2h.
      • Test high-rate performance (~100 kHz) 2h.
      • Commissioning triggers 2h/days
        • Assess trigger efficiencies
        • Set-up specific triggers TagH/TagM & ST for PS acceptance and efficiency studies, TAC, ST&FCal&TOF (track enriched trigger) for FDC calibration
          • Once Tagger trigger is established, do HV scans PS coarse counters (1-2h, can be done concurrently with other tasks).
            • Need stable beam condition
    • Hall D DAQ tuning. Sergei F. 2h/day for about 4 days.
    • π0 calibration.
      • Need an empty target run early on during the π0 calibration periods. (5h: 1h to empty/fill back the target, 4h running). Spend 2h at the highest current possible and 2h at the same current as for production.
      • For BCal, 25h of data taking. Need 65M events in mode 8 (8h at 3kHz) and 585M in mode 7 (17h at 10kHz). Concurrently FCal is taking data too). Done at nominal run condition, 10-4 RL 200nA, if possible.
      • Run overnight. Part of trigger optimization.
    • Goniometer commissioning. (Ken L., Hovanes, Paul) Ken Livingston will be at JLab between Feb. 18th-27th. Expected time: 12h in 4 blocks for the 50 μm diamond (if things go well). No plan to work on other thick diamonds unless something is wrong with the 50 μm diamond (damaged, etc.). To be scheduled on morning when Ken is on site, night task otherwise.
    • Trigger Level-3 commissioning. 4h/day for 2 days early in the run. Then 4h at the end of the run. Data need to be taken between production data for comparison.
    • Run PS with different magnetic field (Alex S., 0h: done in parallel)
  4. Hall D run
    • Continue above Hall D tasks if needed, then:
      • TOF HV scan (3h, Alex. Ostrovidov). Plan for TOF HV scan.
        • Drift chambers in their optimal state and calibrated.
        • π0 calibration data can be taken concurrently.
      • Microscope alignment (Alex B.) Iterative process:4×20 min, spread throughout the run.
        • Require stable beam of known position (from harp scans) ⇒ to be done after FFB commissioning
    • Establish absolute photon flux with PS and TAC (Hovanes E., Alex S., (1+2+2?)h×3 + 4h×3)
      • Prerequisite: nA BPM commissioning preferable but not necessary, since Accelerator has been able to run at a few nA before.
      • Target should be empty (30min to empty, 30 min to refill).
      • Perhaps install a plug in the 5mm collimator hole to decrease photon flux (1h)? Need to de-install the plug once test is done (1h). Need RadCon presence for work on the beamline (ARM not sufficient).
      • Iterative procedure: 2h/days for 3 days.
      • After successful commissioning of the TAC, do a TAC run every few days (1h run + 1h target empty/fill operation). We will do it 4 times.
    • Physics production data with test diamond (50μm?) and 5 mm hole. We should run at least 12h to have enough data for publication and optimally 0.5 week (Justin S.). This assume 100% efficiency, 1007 γ/s in coherent peak (=220 nA on nominal 20μm production diamond). Ratio of longit./transverse. data:50/50. Switch every 2 runs. Time to switch from longit./parall.: 10min.
      • Make sure PS is running at all times for beam polarization monitoring.
      • Need unpolarized data interspersed between pol. run. 20% of the pol. data at least at the beginning. Time to switch from pol. to unpol.: 10min. time to switch from unpol. to pol.: 30min.
      • Before the end of this task, take a 4h empty target run with polarized beam. Can be done at the same time as a TAC run, to save 1h on emptying/filling the target.
    • Test effect of quadrupole on microscope data. Alex B./Alex. S./Todd S. Iterative work: 4h/days for 4 days over a week. Production data can be taken concurrently, but quad. test has priority. Need also to scan magnetic to find optimal configuration. Alex S. calculation indicates it is 20% above nominal set current of 3.2968A.
    • Thin diamond test and production (characterized thin diamonds will be available early March)
      • Install thin diamonds (Install between early and mid-March. 8h but opportunistic: done during beam down time).
      • Establish polarized beam with on the 3 thin diamonds. (Hovanes, Paul) 3 days
      • If polarized beam is succesfuly established, then:
        • Take data with thin diamond and 5 mm hole . Expected time (Justin S.), same as for 50μm diamond run, counting these 50μm diamond data.
          • Before the end of this task, take a 4h empty target run with polarized beam. Can be done at the same time as a TAC run, to save 1h on emptying/filling the target.
        • Move A.C. to 3.4 mm hole and check beam characteristics with 3.4 mm hole. 1h Need RadCon presence for work on the beamline (ARM not sufficient).
        • Take data with thin diamond and 3.4 mm hole. Expected time (Justin S.): whatever we can take once we have completed what we need on 5mm collimator.
          • Before the end of this task, take a 4h empty target run with polarized beam. Can be done at the same time as a TAC run, to save 1h on emptying/filling the target.
    • opportunistic TRD test.
    • opportunistic polarimeter test.

Hall D white board (current run plan)

Hall D white board 1:
Hdwb1.JPG
Hall D white board 2:
Hdwb2.JPG

Runplan time charts (Original plan)

Commissioning time chart: week 1:
Week1.jpg

Commissioning time chart. Time t=0 corresponds to 02/05/2016. The horizontal scale is the time during which beam is available (in other word, it would be the real time with 100% of beam efficiency). The task time estimates given on the plot have 100% contingency.

Commissioning time chart: week 2:
Week2.jpg
Commissioning time chart: week 3:
Week3.jpg
'Commissioning time chart: week 4:
Week4.jpg
Commissioning time chart: week 5:
Week5.jpg
Commissioning time chart: week 6:
Week6.jpg
Commissioning time chart: week 7:
Week7.jpg
Commissioning time chart: week 8:
Week8.jpg
Commissioning time chart: week 9:
Week9.jpg

Runplan time charts (readjusted after CHL repair started on March 9th)

Readjusted commissioning time chart: week 8:
Commissioning time chart adjusted for the CHL repairs. Time t=0 corresponds to 03/27/2016. The horizontal scale is the time during which beam is available (in other word, it would be the real time with 100% of beam efficiency). The task time estimates given on the plot have 100% contingency.
Readjusted commissioning time chart: week 9:
Week9 3.jpg
Readjusted commissioning time chart: week 10:
Week10 3.jpg
Readjusted commissioning time chart: week 11:
Week11 3.jpg

Accelerator status

Hot Checkout

  1. Hot Check OUT Status 100% ready ; 0% checked ; 0% not ready

Discussion on performances and problems during the Fall 15 run. Subsystem preparations and goals for the Spring 16 run

Summary of the Fall 15 run, with details on subsystem performances

  • Solenoid (Eugene/Beni)
  • DAQSurveys
    • Polarimeters
  • DAQ (Sergei)
    • DAQ is ready.
  • Trigger (Alex S.)
  • Beam line
    • Collimator plug (Richard J./Tim). Goal: reduce photon flux by factor ~10. Proposed characteristics: Tungsten, 7.5 cm long, 1.05mm ID (.041 inches) +/- 0.05mm. Timeframe to machine: about 2 weeks. Document summarizing the simulation study from Richard. J.
    • Tagger hodoscope (Nathan/Franz)
      • Obtained good data from voltage scan done during Fall 15 run. Can now use lower UV.
      • As of Jan. 13th, 5 counters do not work. They will be replaced on Feb. 1st.
    • Tagger microscope (Alex B./Richard)
      • Fall 15 data allowed to optimized bias voltage. Know thresholds of ADC and discrim. As of of Jan. 13th, need to re-solder 5-6 SiPM. Otherwise, it is ready.
    • Tagger quadrupole tested. Other magnets tested as well.
    • Goniometer (Hovanes)
    • Active collimator, profiler, flux monitor (Alex B./Hovanes)
    • PS system (Alex S.)
      • 5 channels are still not working. The will be fixed after the end of the Spring 16 run.
      • Threshold adjustment needed for high granularity counter.
      • Need to calibrate Hall probe (done on Jan. 14th)
    • TRD (Sergei, Lubomir)
      • Small leak on gas line. Will fix it before we start.
      • Note for planning: when Xe starts to be used, need to take data (concurrent with main Hall D/GlueX commissioning) within 2-3 days.
    • Polarimeter (M. Dugger/N. Sparks)
      • Adding filtering electronics. It works well.
    • Target (Alexandre D.)
      • Target is ready and pushed in.
    • TAC (Hovanes)
      • Cabled and surveyed.
      • Motion and signals checked.
      • Need to wait for tagger trigger to work well before to make TAC plans.
    • Slow control (Hovanes)
      • Slow controls are ready
  • Detector systems
    • ST (Mahmoud)
      • ST is ready. HCO done.
      • Tim is designing a mechanism to avoid HV cables to be pulled off when the target is moved.
      • When bias voltage are changed, all other running condition should be maintained to insure meaningful comparison.
    • FDC (Lubomir)
      • FDC are ready.
      • Some issues with HV channels related to low Hall humidity
      • No straight track data necessary, unless an opportunity arises. Will then take 4h of data.
    • CDC (Beni). Ready
    • FCal (Adesh)
      • May need to replace 2-3 bases.
      • EPICS GUI not working -> cannot test FCal. Fixed now. Able to take pulser data.
      • 12 channels have HV problems. 20 channels were not reported because they were not reconnected or FADC not working or PMT base not working.
    • Bcal (Elton)
      • Bcal chiller temperature was lower to 5 degree C.
    • TOF (Alex. Ostrovidov) cosmic run done. Ready for run.
  • Online monitoring (David L.)
  • Trigger level 3 (David L.)


  • Physics data taking (Justin S.)


  • RC duty:
    • Thur. February 4th to Wed. March 14th: Alexandre Deur
    • Wed. March 15th - Fri. March 29th: Mark Ito
    • Fri. March 30th - Wed. April 13th: David Lawrence
  • Other topics

Meeting Notes